Two years ago, Bethesda Corporation bought a delivery truck for $30,000 (not subject to the luxury auto depreciation limits). Bethesda used MACRS 200 percent declining balance and the half-year convention to recover the cost of the truck, but it did not elect §179 expensing or eligible bonus depreciation. Answer the questions for the following alternative scenarios.

a. Assuming Bethesda used the truck until it sold it in March of year 3, what depreciation expense can it claim on the truck for years 1 through 3?

b. Assume that Bethesda claimed $18,500 of depreciation expense on the truck before it sold it in year 3. What is the amount and character of the gain or loss if Bethesda sold the truck in year 3 for $17,000, and incurred $2,000 of selling expenses on the sale?

c. Assume that Bethesda claimed $18,500 of depreciation expense on the truck before it sold it in year 3. What is the amount and character of the gain or loss if Bethesda sold the truck in year 3 for $35,000, and incurred $3,000 of selling expenses on the sale?

d. Assume that Bethesda claimed $18,500 of depreciation expense on the truck before it sold it in year 3. What is the amount and character of the gain or loss if Bethesda sold the truck in year 3 for $7,000, and incurred $2,000 of selling expenses on the sale?
Answer #1

Ansu Jear- Depreialionongina basii x Depvecation e 3 0000X O.2 600 0 ea11 Depreciali on - 30000 X O 32 9600 year 2 Pepretiatisoda ddeivengtuk lacs Compuet gain ding pens Amount cali2ed price 1000-2000 SO00 = 30000 - 18500 Adjuckd basis 1sO0 GoinClocTCompute gails on ale d delivey tutk Amou nt alid 35000-3co0 Adjuntd basis 30000 185O0 |\500 Giainosstakad- 32000 --1500 Parti2ddutiuy tuck On 000-aooo Amourt ealiaed S000 ad jued bacs 30000-8500 Gainloss)ta iad = 11500 - 5000 GS00 Amourt Amouri) Part